Bob is three times older than Alice. Ten years ago Bob's age was twice that what Alice's age will be in 5 years. How old are Bob and Alice now?
Answer by Cromlix(4381)

Bob is three times older than Alice.
B = 3A
Ten years ago Bob's age was twice that what Alice's age will be in 5 years.
B - 10 = 2(A + 5)
Replace B with 3A from first equation.
3A - 10 = 2(A + 5)
3A - 10 = 2A + 10
Collect like terms
3A - 2A = 10 + 10
A = 20
Currently;
Alice is 20 years old
Bob is 60 years old.
Hope this helps.
:-)